OpenStudy (naveenbhatia1312):

Find a ⋅ b. a = <7, 4>, b = <3, 5> 41 <21, 20> <10, 9> -1

Directrix (directrix):

What are the instructions for this problem?

OpenStudy (naveenbhatia1312):

find a * b so i got 41. is that correct

Directrix (directrix):

For the dot product, a*b = 7 · 3 + 4 · 5 = 21 + 20 = 41 is what I got which is the same as yours.

OpenStudy (naveenbhatia1312):

can i ask 1 more question

OpenStudy (naveenbhatia1312):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8> 3.0° 6.0° 16.0° -7.0°

Directrix (directrix):

On the second part of the first problem, <21, 20>* <10, 9>, I did not get -1 for the dot product. Check your work on that. @naveenbhatia1312

OpenStudy (naveenbhatia1312):

i did not either. I got 41

Directrix (directrix):

Okay, so the first problem is just one problem, then. Hold on while I look for the angle.

Directrix (directrix):

cos(x) = .994505 x = 6.0° --> Did you get this?

OpenStudy (naveenbhatia1312):

yes

OpenStudy (naveenbhatia1312):

thanks

Directrix (directrix):

You are welcome.